OpenStudy (anonymous):

a particle moving in a straight line is given by the equation of motion s=1/t^2. find the velocity of the particle at t=a

OpenStudy (anonymous):

Using \[\lim_{h \rightarrow 0} \frac{ F(a +h) -F(a) }{ h }\]

OpenStudy (anonymous):

It doesn't look like a particularly hard limit

OpenStudy (whpalmer4):

You want to find the derivative (by first principle, apparently) of s(t) because v = ds/dt. \[\frac{ds}{dt} = \lim_{x \rightarrow 0}{\frac{1}{h}(\frac{1}{(t+h)^2} - \frac{1}{t^2} )}\] Once you've ground that out, plug in t = a.

OpenStudy (whpalmer4):

Sorry, that limit is obviously as h->0, not x!

OpenStudy (anonymous):

what happened to the H in the denominator?

OpenStudy (whpalmer4):

I wrote it as (1/h) times the numerator

OpenStudy (anonymous):

ok I am still a bit confused because I thought is was times the reciprocal so it would be h/1?

OpenStudy (anonymous):

brain food is done be back in 2 min

OpenStudy (whpalmer4):

No, the messy thing in the parentheses is just the numerator of the original fraction. Here, I'll write it out in the original form, so you can see why I don't do that! \[\frac{ds}{dt} = \lim_{h\rightarrow 0} \frac{\frac{1}{(t+h)^2}-\frac{1}{t^2}}{h}\]

OpenStudy (whpalmer4):

I think it's easier to deal with it written like this: \[\frac{ds}{dt} = \lim_{h \rightarrow 0} (\frac{1}{h})(\frac{1}{(t+h)^2} - \frac{1}{t^2} )\] After you multiply both top and bottom of the left fraction in the parentheses by t^2 and the top and bottom of the right fraction by (t+h)^2 you can factor an h out in the numerator to cancel the 1/h, and then just take the limit.

OpenStudy (anonymous):

@adam32885 Is the problem that you don't know how to subtract fractions with variables in them?

OpenStudy (anonymous):

ok I think i was trying to make things harder than it was meant to be let me see what i get

OpenStudy (anonymous):

Okay I am still messing up somewhere

OpenStudy (anonymous):

Do you know how to subtract fractions?

OpenStudy (anonymous):

yes find a common denominator

OpenStudy (anonymous):

In this case what is it?

OpenStudy (anonymous):

\[t ^{2}(t+h)^{2}\]

OpenStudy (anonymous):

okay now you have to change the fractions

OpenStudy (anonymous):

so it would be \[\frac{ t ^{2}- (t+h)^{2} }{ t ^{2}(t+h)^{2} }\]

OpenStudy (anonymous):

Yeah that's it. Now simplify

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

No, start by expanding the numerator.

OpenStudy (anonymous):

remember foil?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

actually....

OpenStudy (anonymous):

yeah basically you need to expand the numerator

OpenStudy (anonymous):

just the numerator?

OpenStudy (anonymous):

so I get \[\frac{ -2th+h ^{2} }{ t ^{2}(t+h)^{2} }\]

OpenStudy (anonymous):

yeah just the numerator for now

OpenStudy (anonymous):

hurry, I gotta go soon.

OpenStudy (anonymous):

go go go

OpenStudy (anonymous):

Now you can divide by \(h\)

OpenStudy (anonymous):

and evaluate the limit

OpenStudy (anonymous):

I am still coming up with a 0 in the numerator

OpenStudy (anonymous):

hell no.

OpenStudy (anonymous):

\[ t^2(t+0)^2 \neq 0 \]

OpenStudy (anonymous):

I was just about to say i see what i did thank you very much

OpenStudy (anonymous):

\[ h-2t \neq 0 \]

OpenStudy (anonymous):

done?

OpenStudy (anonymous):

yes thank you